The width of a rectangle is 10 m less then half the length. If the perimeter is 52 meters, find the length of each side of the rectangle.
Answer by stanbon(75887)

The width of a rectangle is 10 m less then half the length. If the perimeter is 52 meters, find the length of each side of the rectangle.
------
Let the length be "L":
Then the width = (1/2)L-10
------
Perimeter = 2(length + width)
52 = 2(L + (1/2)L-10)
26 = (3/2)L - 10
(3/2)L = 36
Length = 24 meters
-------
Width = (1/2)L-10 = 2 meters
-------------------------------
Cheers,
Stan H.
-------------------------------